Money and the Money Market Using the money market graph, show the effects of an increase...

  1. Money and the Money Market
    1. Using the money market graph, show the effects of an increase in the price level. Please describe the changes beside your graph. (5 marks)
    2. In an economy for the month of June 2019, individuals and businesses held $35 billion in currency, there was $246 billion in personal chequable deposits and $325 billion in non-personal chequable deposits; non-chequable personal deposits were $164 billion and non-chequable business deposits were $39 billion, and term deposits totaled $314 billion.

  1. Calculate the values for M1. (2.5 marks)
  2. Calculate the value for M2. (2.5 marks)
Answer #1

(a)

An increase in price level lowers purchasing power, which increases the demand for money. The money demand curve shifts rightward, increasing interest rate.

In following graph, MD0 and MS0 are initial demand and supply curves for money, intersecting at point A with initial interest rate r0 and quantity of money M0. As money demand rises, MD0 shifts right to MD1, intersecting MS0 at point B with higher interest rate r1 and same quantity of money M0.

(b)

(i) M1 ($ billion) = Currency + Personal checkable deposits + Non-Personal checkable deposits

= 35 + 246 + 325

= 606

(ii) M2 ($ billion) = M1 + Non-checkable personal deposits + Non-chckuable business deposits + Term deposits

= 606 + 164 + 39 + 314

= 1,123